When does an investor make a profit given the following information?

Buy European Put for 3
Stock Price is Currently 42
Stock Price is 40
Answer #1

Stock Price = S = $ 42, Strike = $ 40 and Put Premium (Price) = $ 3

A put option is in the money only when the stock price is below the strike price of the put. Hence, the stock price has to go below $ 40 for the put to be in profit.

If the stock price is $ S, then Net Profit = (40 - S) - Put Premium = (40-S) - 3 = 37 - S

Hence, the position is profitable only when the stock price goes below $ 37.